find the ordered pair
-2 x +5 y=-30
a) (-10,_)
b) (_,-3)
c) (_,8)

Answer by richwmiller(17219)

one after another
plug in the -10 for x and find y
-2(-10)+5y=-30
plug in -3 for y and find x
-2x+5(-3)=-30
plug in 8 for y and find x
-2x+5(8)=-30
reminder:no similar problems